The probability that? Tope passes an exam is 1/3 and probabilities that his brother and. sister pass the same exam are 2/5 and 3/8 respectively. Find the probability that the three of them will fail the exam

a. 0.3333
b. 0.25
c. 0.375
d. 0.4
1
Expert's answer
2020-12-23T16:55:58-0500

Tope fail,P=1-1/3=2/3

brother fail, P=1-2/5=3/5

sister fail, P=1-3/8=5/8

 three of them will fail the exam:

P=(2/3)*(3/5)*(5/8)=0.25